# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ 4.3. CR characterization of Weyl metrizability
Now we can use Proposition 4.4 to show that the necessary conditions in Proposition 4.1 are also sufficient. We first give a Cartan geometric proof of the following well-known lemma, e.g. see [LeB80, Met20], which shows that given a projective structure [],delimited-[][$\nabla$],[ ∇ ] , any conformal structure $[g]$ can have at most one conformal connection that satisfies $\nabla^{[g]}\in[\nabla]$.
@@
If a projective structure []delimited-[][$\nabla$][ ∇ ] is Weyl metrizable with respect to a conformal structure $[g],$ then the property $\nabla^{[g]}\in[\nabla]$ determines a unique conformal connection for $[g].$
@@
holds where, in relation to (2.1), one has $[\omega^{i}_{j}]\in\Omega^{1}(\mathcal{C}_{0},\mathfrak{co}(p,q)),p+q=3,$ and $\omega^{1}_{0}=s_{1}^{*}\theta^{1},\omega^{2}_{0}=s_{1}^{*}\theta^{2},\omega^{1}_{2}=s_{1}^{*}\theta^{3},\omega^{0}_{0}=s_{1}^{*}\theta^{0}.$ As was mentioned in § 2.1, any other conformal connection is related to $s_{1}$ by an action of $P_{+}$ at each point. In other words, if $s_{2}\colon\mathcal{C}_{0}\to\mathcal{C}$ is another conformal connection, then at every $p\in\mathcal{C}_{0},$ $s_{2}(p)=\mathbf{g}^{-1}s_{1}(p),$ for some $\mathbf{g}\in P_{+},$ i.e.
@@
where $\psi_{i}=s_{i}^{*}\psi.$ Explicitly, expressing elements of $P_{+}$ as $\mathbf{g}\in\mathrm{SO}(p+1,q+1)$ where
$\mathbf{g}=\begin{pmatrix}1&q_{2}&q_{1}&q_{0}&\textstyle{\frac{1}{2}}\varepsilon_{ij}q^{i}q^{j}\\ 0&1&0&0&\varepsilon q_{2}\\ 0&0&1&0&\varepsilon q_{1}\\ 0&0&0&1&q_{0}\\ 0&0&0&0&1\\ \end{pmatrix},$
@@
Carrying out the same steps in order to relate two projective connections in a projective structure []delimited-[][$\nabla$][ ∇ ] leads to (3.1), from which it follows that two conformal connections $\nabla^{1}$ and $\nabla^{2}$ are projectively equivalent if and only if their respective connection forms $\omega^{i}_{j}$ and $R^{*}_{q}\omega^{i}_{j}$ are related by
@@
Summing over $i,j,$ one obtains $r_{k}=\tfrac{3}{4}q_{k}.$ Subsequent insertion into (4.16) and lowering the indices using $\varepsilon_{ij}$, one can skew-symmetrize both sides to obtain $q_{0}=q_{1}=q_{2}=0.$
@@
Given an oriented projective structure []delimited-[][$\nabla$][ ∇ ] on a 3-manifold $M,$ there is a one-to-one correspondence between conformal structures $[g]$ with respect to which []delimited-[][$\nabla$][ ∇ ] is Weyl metrizable and Levi nondegenerate CR submanifolds of hypersurface type that are transverse to the Levi kernel of $\mathcal{T}^{[\nabla]}$ whose fundamental binary quartic is either zero or has a repeated root of multiplicity four and satisfy (4.4) for $\varepsilon=1$ or $-1$ with respect to the canonical CR connection described in Proposition 4.4
@@
By Proposition 4.1, we only need to prove the sufficiency part. By the proof of Proposition 4.4, a holomorphic section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M,$ defines a nondegenerate CR structure on $N$ with a canonical CR connection whose fundamental binary quartic has coefficients (4.12).
@@
$\tfrac{\partial}{\partial\overline{\zeta}^{1}}\Im Y_{2}-\tfrac{\partial}{\partial\overline{\zeta}^{2}}\Im Y_{1}=$ $-\tfrac{3}{4}\mathrm{i}Y_{0},$
@@
The condition that the fundamental binary quartic has a repeated root of multiplicity four implies
@@
in (4.12). It follows from (4.18), together with $\Re Y_{1}=0$ from (4.11), that
@@
$3X_{0}=$ $2W^{0}_{221}\Re Y_{2}-5W^{2}_{120}-W^{1}_{202}-4W^{2}_{201}+\mathrm{i}\left(2W^{0}_{112}\Re Y_{2}+5W^{2}_{220}-W^{2}_{101}-4W^{1}_{102}\right).$
The differential consequences of the relations above are straightforward to obtain and we will only provide the final result as algebraic relations among the entries of $W^{i}_{jkl}$ and $W_{ijk}$ in (3.3):
@@
Lastly, using $\mathrm{d}^{2}\tau^{1}=\mathrm{d}^{2}\tau^{2}=0,$ it is straightforward to obtain $\mathrm{d}\Re Y_{2}=2\Re Y_{2}(\eta^{0}_{0}-\eta^{1}_{1})+\Re Y_{2;0}\omega^{0}.$ Moreover, by Bianchi identities one has $\Re Y_{0;3}=\Re Y_{2;0}.$ Using $Y_{0}=0$ in (4.19), it follows that $\Re Y_{2}$ is a constant up to positive homotheties, i.e.
@@
Since by (4.17) it is assumed that $Y_{2}\neq 0,$ there is a natural sub-bundle $\iota_{4}\colon\mathcal{P}_{4}\to\mathcal{P}_{3}$ defined as
$\mathcal{P}_{4}=\left\{p\in\mathcal{P}_{3}\ \vline\ \Re Y_{2}(p)=-\varepsilon\right\}$
where $\varepsilon$ can be $+1$ or $-1$ and (4.17) takes the form (4.4) since $\Im Y_{2}=0$ by (4.19). As a result, $\mathcal{P}_{4}\to N$ is a principal $\mathrm{CO}(2,\mathbb{R})$-bundle. Defining $\iota:=\iota_{1}\circ\iota_{2}\circ\iota_{3}\circ\iota_{4}\colon\mathcal{P}_{4}\to\mathcal{P},$ one obtains
@@
where $\varphi$ is given in (3.6) and
@@
By the homogeneity of $\iota^{*}\varphi$ in (4.20) when pulled-back to the leaves of $\{\tau^{0},\tau^{1},\tau^{2}\},$ it is clear that the (open) fibers of $\mathcal{P}_{4}\to M,$ can be completed to $\mathrm{CO}^{o}(p,q).$ Since the fibers of $\mathcal{P}_{4}\to N,N\subset\mathbb{P}_{+}T^{*}M,$ are $\mathrm{CO}(2,\mathbb{R}),$ it follows that the (open) fibers of $N\to M$ can be completed to $\mathrm{CO}^{o}(p,q)/\mathrm{CO}(2,\mathbb{R})$ as discussed in § 2.4, where $\varepsilon=1$ or $-1$ if $(p,q)=(3,0)$ or $(2,1)$, respectively. ∎
@@
Given an oriented projective structure []delimited-[][$\nabla$][ ∇ ] on a 3-manifold $M,$ suppose $N\subset\mathbb{P}_{+}T^{*}M$ is a Levi nondegenerate CR submanifold of hypersurface type that is transverse to the Levi kernel of $\mathcal{T}^{[\nabla]}$ whose fundamental binary quartic is either zero or has a repeated root of multiplicity. Then either $N$ corresponds to a conformal connection or it satisfies (4.4) for $\varepsilon=0$ in which case $\mathcal{T}^{[\nabla]}$ is locally fibered over a complex curve and a distinguished line field is induced on $M.$ Assuming analyticity, the local generality of such nondegenerate CR 5-manifolds for which $\varepsilon=0$ is 4 functions of 3 variables.
@@
By Theorem 4.7, when $\varepsilon=\pm 1$ the nondegenerate CR structure on $N$ corresponds to a conformal connection. When $\varepsilon=0,$ then, by (4.4), the Pfaffian system $\mathcal{J}=\{\zeta^{1},\overline{\zeta}^{1}\}=\{\tau^{1},\tau^{2}\}$ is integrable. The leaf space of the 3-dimensional leaves of $\mathcal{J}$ is a complex curve equipped with the coframing $(\zeta^{1},\overline{\zeta}^{1}).$ One can check that the Pfaffian system $\mathcal{J}$ is well-defined on $M$ since it is preserved by the action of the fibers of $\pi_{3}\colon\mathcal{P}_{3}\to M$ as defined in (4.11) i.e. for any vertical vector field $X$ in $\pi_{3}:\mathcal{P}_{3}\to M$ the Lie derivative of $\mathcal{J}$ satisfies $\mathcal{L}_{X}\mathcal{J}=\mathcal{J}.$ Thus, $\mathrm{Ker}\mathcal{J}$ defines a line field $\ell\subset TM$ where $\ell=\langle(\pi_{3})_{*}\tfrac{\partial}{\partial\tau^{0}}\rangle.$ In fact, one has $\mathrm{Ker}\mathcal{J}=\ell\oplus\mathscr{V}$ where $\mathscr{V}$ is the rank 2 vertical distribution for $N\to M.$ Carrying out the computations to find the resulting structure equations and algebraic relations among the components of the projective Weyl curvature, which are tedious but straightforward, goes similarly to that of Theorem 4.7 with the difference that in (4.19) in general one does not have $a_{0}=0.$ We will not provide the details here. Doing so will allow one to use Cartan-Kähler machinery and obtain the local generality of such CR structures to be 4 functions of 3 variables. ∎
@@
The class of Levi nondegenerate CR structures in Corollary 4.8 satisfy (4.4) for $\varepsilon=1,-1$ or $0$ If $\varepsilon=\pm 1,$ then they are foliated by null Lagrangian surfaces and correspond to conformal connections and, by our discussion in § 1, their local generality depends on 5 functions of 3 variables. If $\varepsilon=0,$ then they are locally foliated by 3-dimensional leaves. One can analyze the structure of the 3-dimensional leaves of the latter class of CR structures. To do so, one sets $\tau^{1}=\tau^{2}=0$ in the structure equations and obtains that the 1-form $\tau^{0}$ is integrable on each leaf. The respective leaves of $\{\tau^{0},\tau^{1},\tau^{2}\}$ were discussed in § 2.3 as holomorphic curves that are null and Legendrian, c.f. [Bry82]. Inline with the result of [LeB85] on foliated CR manifolds with compact complex manifolds, one can ask whether there is an example for which all such 3-dimensional leaves are compact.
@@ 5.1. Conformal rigidity of flat projective structures
A 3-dimensional conformal structure has a projectively flat conformal connection if and only if it is flat.
@@
Alternatively, one can give a more direct proof of the first part as follows. Given a conformal connection $\nabla^{{[g]}},$ associated to $s\colon\mathcal{C}_{0}\to\mathcal{C},$ one has (4.1) and (4.2), using which the projective Weyl curvature of $[\nabla^{[g]}]$ can be expressed in terms of $\mathsf{P},$ e.g.
(5.1) $\begin{gathered}W^{0}_{112}=\varepsilon\mathsf{P}_{02},\quad W^{0}_{212}=-\varepsilon\mathsf{P}_{01},\quad W^{1}_{002}=\varepsilon\mathsf{P}_{12},\quad W^{1}_{202}=-\mathsf{P}_{10},\\ W^{1}_{212}=\textstyle{\frac{1}{2}}(\varepsilon\mathsf{P}_{00}-\mathsf{P}_{11}),\quad W^{2}_{001}=\varepsilon\mathsf{P}_{21},\quad W^{2}_{101}=-\mathsf{P}_{20},\quad W^{2}_{121}=\textstyle{\frac{1}{2}}(\varepsilon\mathsf{P}_{00}-\mathsf{P}_{22}),\\ \end{gathered}$
where $\varepsilon=1$ or $-1$ in the Riemannian or Lorentzian signature, respectively. Projective flatness is equivalent to $W^{i}_{jkl}=0,$ which by (5.1) implies
@@
By Remark 4.2 one obtains that $\nabla^{[g]}$ is locally the Levi-Civita connection of an Einstein metric in $[g]$ and therefore $\mathsf{P}_{00}$ is a constant defined up to positive homotheties on $\mathcal{C}_{0}.$ Moreover, the entries of the Cotton-York tensor, denoted as $Y_{ijk}$ in (2.4), are determined by the first jet of the Rho tensor $\mathsf{P}_{ij}$ for any conformal connection via the relation
@@
A conformal structure in dimension $\geq 3$ has a projectively flat conformal connection if and only if it is flat.
@@
and $[\varepsilon_{ij}]$ is the standard diagonal matrix of signature $(p,q)$ and $p+q=n.$ By our discussion in § 2.2, a (pseudo-)conformal structure of signature $(p,q)$ on an $n$-dimensional manifold $M$ corresponds to a Cartan geometry $(\mathcal{C}\to M,\psi)$ of type $(\mathrm{SO}(p+1,q+1),Q_{1}).$ The structure equations on $\mathcal{C}$ is known to be, e.g. see [Kob72, Chapter 4],
@@
The conformal structure on $M$ is $[t^{*}g]\subset\mathrm{Sym}^{2}(T^{*}M),$ for any section $t\colon M\to\mathcal{C}$, where $g=\varepsilon_{ij}\omega^{i}\omega^{j}\in\Gamma(\mathrm{Sym^{2}}(T^{*}\mathcal{C}))$ Furthermore, $[\theta^{i}_{j}]$ is $\mathfrak{co}(p,q)$-valued and the Weyl curvature entries $C^{i}_{jkl}$ satisfy
@@
Similarly, by our discussion in § 3.1, a projective structure []delimited-[][$\nabla$][ ∇ ] on an $n$-dimensional manifold is a Cartan geometry $(\mathcal{P}\to M,\varphi)$ of type $(\mathrm{SL}(n+1,\mathbb{R}),P_{1})$ for which the structure equations are, e.g. see [Kob72, Chapter 4],
@@
where $[\omega^{i}_{j}]$ is $\mathfrak{gl}(n)$-valued. We point out that the 1-forms $\omega^{i}_{j}$ above and $\eta^{i}_{j}$ in (3.2) are related by $\omega^{i}_{j}=\eta^{i}_{j}+\delta^{i}_{j}\eta^{k}_{k}.$ The entries of the projective Weyl curvature $W^{i}_{jkl}$ satisfy
@@
Viewing the conformal connection $\nabla^{[g]}$ as a projective connection for $[\nabla^{[g]}],$ it defines a section $u\colon\mathcal{P}_{0}\to\mathcal{P},$ where
@@
Since $W_{ij}$ is trace-free, it follows from (5.7) that $W_{ij}=0$ implies $\mathsf{P}_{ij}=c\varepsilon_{ij}$ for some $c,$ which, by Remark 4.2, means $\nabla^{[g]}$ is the Levi-Civita connection of an Einstein metric in $[g].$ Since projective flatness is equivalent to $W_{ijkl}=0,$ one has $\mathsf{P}_{ij}=c\varepsilon_{ij}$ and, consequently, the relation (5.6) gives $C_{ijkl}=0,$ i.e. $[g]$ is a flat conformal structure. As a result, if $[\nabla^{[g]}]$ is projectively flat, then $\nabla^{[g]}$ is the Levi-Civita connection of (pseudo-)Riemannian metrics of constant sectional curvature.
@@
Although projectively flat Weyl structures are rigid in dimension $\geq 3,$ one can show, e.g. using Cartan-Kähler theory, that on surfaces they locally depend on 4 functions of 1 variable. A description of conformal connections whose geodesics are great circles on 2-sphere can be found in [Met14, Met21].

### v3

@@ 4.3. CR characterization of Weyl metrizability
Now we can use Proposition 4.4 to show that the necessary conditions in Proposition 4.1 are also sufficient. We first give a Cartan geometric proof of the following well-known lemma, e.g. see [LeB80, Met20], which shows that given a projective structure $\nabla$], any conformal structure $[g]$ can have at most one conformal connection that satisfies $\nabla^{[g]}\in[\nabla]$.
@@
If a projective structure $\nabla$] is Weyl metrizable with respect to a conformal structure $[g],$ then the property $\nabla^{[g]}\in[\nabla]$ determines a unique conformal connection for $[g].$
@@
holds where, in relation to (2.1), one has $\omega^{i}_{j}]\in\Omega^{1}(\mathcal{C}_{0},\mathfrak{co}(p,q)),p+q=3,$ and the only non-zero 1-form entries in $\omega^{i}_{j}]$ are
$\omega^{1}_{0}=s_{1}^{*}\theta^{1},\qquad\omega^{2}_{0}=s_{1}^{*}\theta^{2},\qquad\omega^{1}_{2}=s_{1}^{*}\theta^{3},\qquad\omega^{0}_{0}=s_{1}^{*}\theta^{0}.$
By Definition 2.2 for a general Weyl structure, any other conformal connection is related to $s_{1}$ by an action of the nilradical factor $Q_{+}\subset Q_{1}$ at each point, where $Q_{1}\subset\mathrm{SO}(p+1,q+1)$ is the first parabolic subgroup, discussed in § 2.2. In other words, if $s_{2}\colon\mathcal{C}_{0}\to\mathcal{C}$ is another conformal connection, then at every $p\in\mathcal{C}_{0},$ $s_{2}(p)=\mathbf{g}_{p}^{-1}s_{1}(p),$ for some $\mathbf{g}_{p}\in Q_{+}\subset Q_{1},$ i.e.
@@
where $\psi_{i}=s_{i}^{*}\psi$ and $\mathbf{g}\in C^{\infty}(\mathcal{C},Q_{+}).$ Explicitly, expressing elements of $Q_{+}\subset\mathrm{SO}(p+1,q+1)$ as
$\mathbf{g}_{p}=\begin{pmatrix}1&q_{2}&q_{1}&q_{0}&\textstyle{\frac{1}{2}}\varepsilon_{ij}q^{i}q^{j}\\ 0&1&0&0&\varepsilon q_{2}\\ 0&0&1&0&\varepsilon q_{1}\\ 0&0&0&1&q_{0}\\ 0&0&0&0&1\\ \end{pmatrix}$,
@@
Carrying out the same steps in order to relate two projective connections in a projective structure $\nabla$] leads to (3.1), from which it follows that two conformal connections $\nabla^{1}$ and $\nabla^{2}$ are projectively equivalent if and only if their respective connection forms $\omega^{i}_{j}$ and $R^{*}_{q}\omega^{i}_{j}$ are related by
@@
Summing over $i,j,$ one obtains $r_{k}=\tfrac{3}{4}q_{k}.$ Subsequent insertion into (4.20) and comparing both sides yields $q_{0}=q_{1}=q_{2}=0.$
@@
Given an oriented projective structure $\nabla$] on a 3-manifold $M,$ there is a one-to-one correspondence between conformal structures $[g]$ with respect to which $\nabla$] is Weyl metrizable and 5-dimensional Levi nondegenerate CR submanifolds of hypersurface type that are transverse to the Levi kernel of the 7-dimensional twistor bundle, $\mathcal{T}^{[\nabla]},$ whose fundamental binary quartic is either zero or has a repeated root of multiplicity four and satisfy the non-integrability condition (4.4) for $\varepsilon=1$ or 1-1 with respect to the canonical CR connection described in Proposition 4.4
@@
By Proposition 4.1, we only need to prove the sufficiency part. By the proof of Proposition 4.4, a holomorphic section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M,$ defines a nondegenerate CR structure on $N$ with a canonical CR connection whose fundamental binary quartic has coefficients (4.16), together with an inclusion $\iota_{012}\colon\mathcal{P}_{2}\to\mathcal{P}$ which is a $\mathrm{CO}(2)\times\mathbb{R}^{+}$-reduction of the Cartan bundle for the projective structure $\nabla$].
@@
(4.22) $\tfrac{\partial}{\partial\overline{\zeta}^{1}}\Im Y_{2}-\tfrac{\partial}{\partial\overline{\zeta}^{2}}\Im Y_{1}=$ $-\tfrac{3}{4}\mathrm{i}Y_{0},$
@@
where $Y_{i}$’s and $X_{i}$’s are the functions obtained from reductions in (4.14) and $W^{i}_{jkl}$’s are the entries of the projective Weyl curvature. The condition that the fundamental binary quartic for the CR structure, (2.9), has a repeated root of multiplicity four implies
@@
Combining the relation in (4.23), (4.16) and (4.22), together with $\Re Y_{1}=0$ from (4.15), it follows that
@@
$3X_{0}=$ $2W^{0}_{221}\Re Y_{2}-5W^{2}_{120}-W^{1}_{202}-4W^{2}_{201}+\mathrm{i}\left(2W^{0}_{112}\Re Y_{2}+5W^{2}_{220}-W^{2}_{101}-4W^{1}_{102}\right),$
holds on $\mathcal{P}_{2}.$ The differential consequences of the relations above are straightforward to obtain and we will only provide the final result as algebraic relations among the entries of $W^{i}_{jkl}$ and $W_{ijk}$ in (3.3):
@@
Lastly, using $\mathrm{d}^{2}\tau^{1}=\mathrm{d}^{2}\tau^{2}=0,$ one obtains
(4.25) $\mathrm{d}\Re Y_{2}=2\Re Y_{2}(\eta^{0}_{0}-\eta^{1}_{1})+\Re Y_{2;0}\omega^{0}.$
Moreover, differential relations between coframe derivatives of $Y_{i}$’s yield $\Re Y_{0;3}=\Re Y_{2;0}.$ Since (4.24) gives $Y_{0}=0$, it follows from (4.25) that $\mathrm{d}\Re Y_{2}=2\Re Y_{2}(\eta^{0}_{0}-\eta^{1}_{1}),$ which means that $\Re Y_{2}$ is a constant up to positive homotheties, i.e.
@@
Since by (4.21) it is assumed that $Y_{2}\neq 0,$ there is a natural sub-bundle $\iota_{3}\colon\mathcal{P}_{3}\to\mathcal{P}_{2}$ defined as
$\mathcal{P}_{3}=\left\{p\in\mathcal{P}_{2}\ \vline\ \Re Y_{2}(p)=-\varepsilon\right\}$
where $\varepsilon$ can be +1+1 or 1-1 and (4.21) takes the form (4.4) since $\Im Y_{2}=0$ by (4.24). As a result, $\mathcal{P}_{3}\to N$ is a principal $\mathrm{CO}(2,\mathbb{R})$-bundle. Defining $\iota:=\iota_{0}\circ\iota_{1}\circ\iota_{2}\circ\iota_{3}\colon\mathcal{P}_{3}\to\mathcal{P},$ one has
@@
which is obtained from the Cartan connection $\varphi$ in (3.6) for the initial projective structure $\nabla$] by applying admissible gauge transformations, assuming (4.23) and $Y_{2}\neq 0$, wherein
@@
By the homogeneity of $\iota^{*}\varphi$ in (4.26) when pulled-back to the leaves of $\{\tau^{0},\tau^{1},\tau^{2}\},$ it is clear that the (open) fibers of $\mathcal{P}_{3}\to M,$ can be completed to $\mathrm{CO}^{o}(p,q).$ Since the fibers of $\mathcal{P}_{3}\to N,N\subset\mathbb{P}_{+}T^{*}M,$ are $\mathrm{CO}(2,\mathbb{R}),$ it follows that the (open) fibers of $N\to M$ can be completed to $\mathrm{CO}^{o}(p,q)/\penalty 50\mathrm{CO}(2,\mathbb{R})$ as discussed in § 2.4, where $\varepsilon=1$ or 1-1 if $(p,q)=(3,0)$ or $(2,1)$, respectively. ∎
@@
Given an oriented projective structure $\nabla$] on a 3-manifold $M,$ suppose $N\subset\mathbb{P}_{+}T^{*}M$ is a Levi nondegenerate CR submanifold of hypersurface type that is transverse to the Levi kernel of $\mathcal{T}^{[\nabla]}$ whose fundamental binary quartic is either zero or has a repeated root of multiplicity. Then either $N$ corresponds to a conformal connection or it satisfies (4.4) for $\varepsilon=0$ in which case $\mathcal{T}^{[\nabla]}$ is locally fibered over a complex curve and a distinguished line field is induced on $M.$ Assuming analyticity, the local generality of such nondegenerate CR 5-manifolds for which $\varepsilon=0$ is 4 functions of 3 variables.
@@
By Theorem 4.7, when $\varepsilon=\pm 1$ the nondegenerate CR structure on $N$ corresponds to a conformal connection. When $\varepsilon=0,$ then, by (4.4), the Pfaffian system $\mathcal{F}=\{\zeta^{1},\overline{\zeta}^{1}\}=\{\tau^{1},\tau^{2}\}$ is integrable. The leaf space of the 3-dimensional leaves of $\mathcal{F}$ is a complex curve equipped with the coframing $(\zeta^{1},\overline{\zeta}^{1}).$ One can check that the Pfaffian system $\mathcal{F}$ is well-defined on $M$ since it is preserved by the action of the fibers of $\pi_{3}\colon\mathcal{P}_{3}\to M$ as defined in (4.15) i.e. for any vertical vector field $X$ in $\pi_{3}:\mathcal{P}_{3}\to M$ the Lie derivative of $\mathcal{F}$ satisfies $\mathcal{L}_{X}\mathcal{F}=\mathcal{F}.$ Thus, $\mathrm{Ker}\mathcal{F}$ defines a line field $\ell\subset TM$ where $\ell=\langle(\pi_{3})_{*}\tfrac{\partial}{\partial\tau^{0}}\rangle.$ In fact, one has $\mathrm{Ker}\mathcal{F}=\ell\oplus\mathscr{V}$ where $\mathscr{V}$ is the rank 2 vertical distribution for $N\to M.$ Carrying out the computations to find the resulting structure equations and algebraic relations among the components of the projective Weyl curvature, which are tedious but straightforward, goes similarly to that of Theorem 4.7 with the difference that in (4.24) in general one does not have $a_{0}=0.$ We will not provide the details here. Doing so will allow one to use Cartan-Kähler machinery and obtain the local generality of such CR structures to be 4 functions of 3 variables. ∎
@@
The class of Levi nondegenerate CR structures in Corollary 4.8 satisfy (4.4) for $\varepsilon=1,-1$ or $0.$ If $\varepsilon=\pm 1,$ then they are foliated by null Lagrangian surfaces and correspond to conformal connections and, by our discussion in § 1, their local generality depends on 5 functions of 3 variables. If $\varepsilon=0,$ then they are locally foliated by 3-dimensional leaves. One can analyze the structure of the 3-dimensional leaves of the latter class of CR structures. To do so, one sets $\tau^{1}=\tau^{2}=0$ in the structure equations and obtains that the 1-form $\tau^{0}$ is integrable on each leaf. The respective leaves of $\{\tau^{0},\tau^{1},\tau^{2}\}$ were discussed in § 2.3 as holomorphic curves that are null and Legendrian, c.f. [Bry82]. Inline with the result of [LeB85] on foliated CR manifolds with compact complex manifolds, one can ask whether there is an example for which all such 3-dimensional leaves are compact.
@@ 5.1. Conformal rigidity of flat projective structures
A 3-dimensional conformal structure locally has a projectively flat conformal connection if and only if it is locally conformally flat.
@@
Alternatively, one can give a more direct proof of the first part as follows. Given a conformal connection $\nabla^{{[g]}},$ associated to $s\colon\mathcal{C}_{0}\to\mathcal{C},$ one has (4.1) and (4.2), using which the projective Weyl curvature of $\nabla^{[g]}]$ can be expressed in terms of $\mathsf{P},$ e.g.
[…]
where $\varepsilon=1$ or 1-1 in the Riemannian or Lorentzian signature, respectively. Projective flatness is equivalent to $W^{i}_{jkl}=0,$ which by (5.1) implies
@@
By Remark 4.2 one obtains that $\nabla^{[g]}$ is locally the Levi-Civita connection of a metric $g_{0}\in[g],$ defined up to positive homotheties, whose Schouten tensor satisfies (5.2) and, therefore, by Schur’s lemma, $\mathsf{P}_{00}$ is a constant defined up to positive homotheties, i.e. $g_{0}$ is Einstein. Moreover, the entries of the Cotton-York tensor, denoted as $Y_{ijk}$ in (2.4), are determined by the first jet of the Rho tensor $\mathsf{P}_{ij}$ for any conformal connection via the relation
@@
A conformal structure in dimension $\geq 3$ locally has a projectively flat conformal connection if and only if it is locally conformally flat.
@@
and $\varepsilon_{ij}]$ is the standard diagonal matrix of signature $(p,q)$ and $p+q=n.$ By our discussion in § 2.2, a (pseudo-)conformal structure of signature $(p,q)$ on an $n$-dimensional manifold $M$ corresponds to a Cartan geometry $(\mathcal{C}\to M,\psi)$ of type $(\mathrm{SO}(p+1,q+1),Q_{1}).$ The structure equations on $\mathcal{C}$ is known to be, e.g. see [Kob72, Chapter 4],
@@
The conformal structure on $M$ is $t^{*}g]\subset\mathrm{Sym}^{2}(T^{*}M),$ for any section $t\colon M\to\mathcal{C}$, where $g=\varepsilon_{ij}\omega^{i}\omega^{j}\in\Gamma(\mathrm{Sym^{2}}(T^{*}\mathcal{C}))$ Furthermore, $\theta^{i}_{j}]$ is $\mathfrak{co}(p,q)$-valued and the Weyl curvature entries $C^{i}_{jkl}$ satisfy
@@
Similarly, by our discussion in § 3.1, a projective structure $\nabla$] on an $n$-dimensional manifold is a Cartan geometry $(\mathcal{P}\to M,\varphi)$ of type $(\mathrm{SL}(n+1,\mathbb{R}),P_{1})$ for which the structure equations are, e.g. see [Kob72, Chapter 4],
@@
where $\omega^{i}_{j}]$ is $\mathfrak{gl}(n)$-valued. We point out that by comparing the expression for $\mathrm{d}\tau^{i}$ using the matrix form (3.2) and using (5.5), it follows that the 1-forms $\omega^{i}_{j}$ above and $\eta^{i}_{j}$ in (3.2) are related by $\omega^{i}_{j}=\eta^{i}_{j}+\delta^{i}_{j}\eta^{k}_{k}.$ The entries of the projective Weyl curvature $W^{i}_{jkl}$ satisfy
@@
Viewing the conformal connection $\nabla^{[g]}$ as a projective connection for $\nabla^{[g]}],$ it defines a section $u\colon\mathcal{P}_{0}\to\mathcal{P},$ where
@@
Since $W_{ij}$ is trace-free, it follows from (5.7) that $W_{ij}=0$ implies $\mathsf{P}_{ij}=c\varepsilon_{ij}$ for some $c,$ which, by Remark 4.2, means $\nabla^{[g]}$ is the Levi-Civita connection of a metric $g\in[g],$ defined up to positive homotheties, whose Schouten tensor satisfies $\mathsf{P}_{ij}=cg_{ij},$ and, therefore, by Schur’s lemma, $c$ is a constant defined up to positive homotheties, i.e. $g$ is Einstein. Since projective flatness is equivalent to $W_{ijkl}=0,$ one has $\mathsf{P}_{ij}=c\varepsilon_{ij}$ and, consequently, the relation (5.6) gives $C_{ijkl}=0,$ i.e. $[g]$ is a flat conformal structure. As a result, if $\nabla^{[g]}]$ is projectively flat, then $\nabla^{[g]}$ is the Levi-Civita connection of (pseudo-)Riemannian metrics of constant sectional curvature.
@@
Although projectively flat Weyl structures are rigid in dimension $\geq 3,$ one can show, e.g. using Cartan-Kähler theory, that on surfaces they locally depend on 2 functions of 1 variable. A description of conformal connections whose geodesics are great circles on 2-sphere can be found in [Met14, Met21].